OpenStudy (anonymous):

write the equation of the line that is parallel to y = 1/2x - 3 and passes through the point (-6,5)

OpenStudy (anonymous):

Hi--- If you recall, in order to write the equation of a line, you need slope and a point on that line

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So, you have the point: (-6, 5)... And, the slope?

OpenStudy (anonymous):

the slope is 1/2 because parallel lines have identical slopes!!

OpenStudy (anonymous):

So, the equation of the line (in point slope form) is y - 5 = 1/2(x + 6)

OpenStudy (anonymous):

ok thanks im having some struggle with algebra 2

OpenStudy (anonymous):

determine whether or not the lines 3x - 4y = 12 and y = -4/3x + 5 are perpendicular. OK... Parallel lines have the same slope..... Perpendicular lines have slopes that are OPPOSITE RECIPROCALS....

OpenStudy (anonymous):

So, what is the slope of each line?